use the discrimant to determine the kinds of numbers that will satisfy the equation x2=3x-4. do not solve.
Answer by mananth(16949)

x^2=3x-4
x^2-3x+4=0
a=1, b=-3,c=4
b^2-4ac
9-4*1*4
=9-16
=-7